OpenStudy (anonymous):

Hi All, could someone help me with this one: 6-3x =5x-10.+18. My instructor gave the answer as being x=6 and I didn't get it all copied down in time today. I cannot figure out how he got that. You group like terms right? Anyway, I'd appreciate any help I might get on this. Denise

OpenStudy (anonymous):

Hi Luis, sorry I was doing something. How did you get that? Could you show me?

OpenStudy (anonymous):

I love how you can use that cool whiteboard thingy?? How do you do that, scanner?

OpenStudy (anonymous):

ohh cool!! Thank you for the info luis:)

OpenStudy (radar):

6-3x=5x-10+18 6-3x=5x+8 -8x=8-6 -8x=2 x=2/-8=- 1/4

OpenStudy (anonymous):

ty Radar, good to see you. Let me take a look at it and try to work it as well. Denise;)

OpenStudy (radar):

I don't see how your instructor got 6??

OpenStudy (anonymous):

Oh it's 10x, sorry, typed it wrong.

OpenStudy (radar):

Is it "6-3x=5x-10x+18" ?

OpenStudy (anonymous):

6-3x=5x-10x+18

OpenStudy (anonymous):

Yes, so sorry radar:( I'm still looking for that lost marble, LOL!

OpenStudy (radar):

Combine similar terms getting: 6-3x=-5x+18 Now add 5x to both sides getting: 6+2x=18 Now sub t ract 6 from both sides getting 2x=12 x=6 Your instrucgtor was correct

OpenStudy (anonymous):

I'm sorry to you too Luis, I typed a period instead of an x after the 10 in the first post:(

OpenStudy (anonymous):

Ok, let me work with that Radar, thank you so much, both of you guys!! Denise

OpenStudy (radar):

For some reason I have lost my delete function key, going to reboot, cul

OpenStudy (anonymous):

Ok. I don't see what happened with the -10x?

OpenStudy (anonymous):

Oh wait, I see, I got it radar. I wasn't lookiing at the combine terms first.

OpenStudy (radar):

Doing a test to determine if my delete key is working. Yea! It is now working.

OpenStudy (radar):

Did you follow with understanding?

OpenStudy (anonymous):

Yes, finally, LOL, it's been awhile since I had to work one like that, now it makes sense, thank you both again:) I can go ahead and do a few more problems now like this one:) See you later and have a nice evening, Denise

OpenStudy (anonymous):

I was trying to combine like terms with terms on the other side of the equal sign

OpenStudy (radar):

Great, good luck in your studies.

OpenStudy (anonymous):

thank you, I surely need it;)

OpenStudy (radar):

Be careful when doing that, it is easy to get the signs twisted.

OpenStudy (anonymous):

that is my biggest problem I think, signs and keeping the rules straight in my head, thank you again.